### v1

# RUIE: Retrieval-based Unified Information Extraction using Large Language Model

Abstract
Unified information extraction (UIE) aims to complete all information extraction tasks using a single model or framework. While previous work has primarily focused on instruction-tuning large language models (LLMs) with constructed datasets, these methods require significant computational resources and struggle to generalize to unseen tasks. To address these limitations, we propose RUIE (Retrieval-based Unified Information Extraction), a framework that leverages in-context learning to enable rapid generalization while reducing computational costs. The key challenge in RUIE is selecting the most beneficial demonstrations for LLMs to effectively handle diverse IE tasks. To achieve this, we integrate LLM preferences for ranking candidate demonstrations and design a keyword-enhanced reward model to capture fine-grained relationships between queries and demonstrations. We then train a bi-encoder retriever for UIE through contrastive learning and knowledge distillation. To the best of our knowledge, RUIE is the first trainable retrieval framework for UIE. Experimental results on 8 held-out datasets demonstrate RUIE’s effectiveness in generalizing to unseen tasks, with average F1-score improvements of 19.22 and 3.13 compared to instruction-tuning methods and other retrievers, respectively. Further analysis confirms RUIE’s adaptability to LLMs of varying sizes and the importance of its key components.
RUIE: Retrieval-based Unified Information Extraction using Large Language Model
Xincheng Liao, Junwen Duanthanks:   Corresponding authors., Yixi Huang, Jianxin Wang Hunan Provincial Key Lab on Bioinformatics, School of Computer Science and Engineering, Central South University, Changsha, Hunan, China {ostars, jwduan, yx.huang}@csu.edu.cn, jxwang@mail.csu.edu.cn
1 Introduction
Figure 1: Illustration of 3 different paradigms for solving unified information extraction task.
Unified Information Extraction (UIE) aims to employ one model or framework to extract diverse structured information from unstructured text, such as Named Entity Recognition (NER), Relationship Extraction (RE), and Event Extraction (EE). Compared to traditional task-specific information extraction (IE) methods, UIE exhibits superior generalization capabilities and higher practical utility Lu et al. (2022); Wang et al. (2023b); Li et al. (2024). Recently, large language models (LLMs) have shown impressive generalization abilities, achieving astonishing results across numerous tasks Wei et al. (2022); Zhang et al. (2024); Wang et al. (2024b). As a result, researchers are leveraging LLMs to tackle the challenges of UIE, which can be categorized into instruction-tuning based method and in-context learning based method (Figure 1).
Due to the significant differences between the structured output of information extraction and large language model pre-training format, current large models still lag behind previous state-of-the-art approaches in IE Han et al. (2023); Ma et al. (2023); Chen et al. (2024). To bridge this gap, some researchers Wang et al. (2023b); Xiao et al. (2024b); Gui et al. (2024); Sainz et al. (2024); Li et al. (2024) have transformed annotated information extraction data into textual or code input-output pairs, training LLMs with instruction-tuning and achieving notable performance. However, instruction-tuning based method faces three main problems: (1) the high training cost, (2) potential loss of general capabilities Xu et al. (2024), (3) poor generalization and performance in unseen tasks.
In addition to instruction-tuning, large language models can also solve problems with just a few examples, which is known as in-context learning Brown et al. (2020). Li et al. (2023); Wang et al. (2023c) have shown that representing structured information in code format, combined with context learning, can outperform small fine-tuned UIE models in few-shot scenarios. Furthermore, Guo et al. (2023) enhances LLM extraction performance by applying offline sentence embedding to encode corpus and searching for relevant examples based on different input tasks. Moreover, some researchers have explored how to obtain more suitable semantic representations for information extraction tasks. Wang et al. (2023a); Wan et al. (2023) use fine-tuned smaller pre-trained language models to encode entities and relations as a retrieval index, achieving performance comparable to previous state-of-the-art methods. Nonetheless, current retrieval-based information extraction methods are mainly tailored to specific IE tasks, lacking UIE generalization ability, and only rely on semantic relevance for examples selection, ignoring the inherent preferences of LLMs.
To make full use of the generalization ability of LLMs and reduce computational costs, we introduce RUIE, a retrieval-based unified information extraction framework. RUIE aims to harness the model’s information extraction capabilities through in-context learning. As shown in Figure 1, compared to instruction tuning-based approaches, RUIE only requires fine-tuning a smaller dense retriever (million-level parameters), greatly reducing the fine-tuning cost. Furthermore, RUIE is compatible with various LLMs, acting as a versatile plugin. Compared to existing retrieval-based information extraction methods, we focus on “true” unified information extraction, where the candidate pool is a mixture of various information extraction tasks such as NER, RE and EE. RUIE is trained to select the most appropriate examples from candidate pool to optimize the LLM’s extraction performance. Furthermore, previous works solely rely on semantic relevance for example selection, we introduce LLM preferences to rank candidate demonstrations. In addition, considering the fine-grained nature of information extraction tasks Wang et al. (2023a); Wan et al. (2023), we have developed a keyword-enhanced reward model to focus on fine-grained information snippets beyond sentence-level semantics. In summary, our main contributions include:
  • We introduce RUIE, the first trainable retrieval framework for Unified Information Extraction (UIE), leveraging in-context learning to reduce computational costs while enabling rapid generalization to unseen tasks.
  • We propose a novel demonstration selection mechanism that incorporates LLM preferences for ranking candidates and utilizes a keyword-enhanced reward model to capture fine-grained relationships between queries and demonstrations.
  • RUIE demonstrates exceptional adaptability, effectively handling diverse IE tasks (NER, RE, EE) and showing compatibility with LLMs of varying sizes, thus serving as a universal plugin for UIE tasks.
2 Related Works
2.1 Unified Information Extraction
Lu et al. (2022) first proposed the framework of unified Information Extraction. They used structured extraction language to unify the input and output forms of IE tasks and the structural schema instructor to guide the model generation, but it requires further fine-tuning for different downstream task. Lou et al. (2023) considers the knowledge transfer between different tasks and schemas, modeling UIE as a semantic matching task with three dimensions of token-token, token-label and lable-token, which achieves better zero/few-shot transfer ability. However, performing three-dimensional semantic matching for each token greatly increases the training and inference cost. Recently, some researchers have modeled IE as natural language Wang et al. (2023b); Gui et al. (2024) or code Sainz et al. (2024); Li et al. (2024) generation tasks. They constructed instruction-tuning datasets based on existing IE datasets to fine-tune large language models, realizing knowledge sharing of different tasks and effectively improving UIE performance. However, the fine-tuning cost of large language models is high, and the fine-tuned models do not generalize well in new domains.
2.2 In-context Learning based Information Extraction
In-context learning (ICL) is the ability of large language models to perform new tasks with only a few examples or demonstrations. One significant merit of ICL is the circumvention of fine-tuning, which might not always be possible due to limited access to the model parameters or constraints on computational resources Brown et al. (2020). Li et al. (2023); Wang et al. (2023c) represent structured IE tasks with codes, and improve extraction performance by randomly selecting examples. Guo et al. (2023) optimizes the random sample retrieval process by introducing sentence embedding to select sample according to the similarity between query and examples. Wang et al. (2023a); Wan et al. (2023) show that fine-grained alignment information such as entities and relations is more important than sentence similarity for example selection in IE tasks. They use the entity and relation representations obtained by the fine-tuned small model to replace the sentence representation for retrieval, and obtain better performance than sentence-level embedding. However, they must obtain the entity span in the sentence and the fine-tuned small model in advance, and the use scenario is limited. To the best of our knowledge, there is no trainable retrieval based framework designed for UIE. On the one hand, we do not need any prior information about the sentence, on the other hand, we design a retrieval training scheme for UIE, which achieves better performance than general sentence embedding.
3 Methods
In this section, we first present the formal definition of information extraction and the task setting for retrieval-based UIE. Then we present our training framework illustrated in Figure 2.
3.1 Problem Statement
IE involves three main tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). For a given sentence $x$, NER seeks to extract tuples $\{s,e\}$, where $s$ represents the entity span and $e$ denotes the entity type. RE focuses on extracting triples $\{e_{s},e_{t},r\}$, with $r$ being the relation type and $e_{s}$ and $e_{t}$ indicating the head and tail entity, respectively. EE comprises two sub-tasks: Event Detection (ED) and Event Argument Extraction (EAE). ED involves extracting event triggers $t\in\mathcal{O}$, where $\mathcal{O}$ is the event type ontology, while EAE extracts arguments $a\in\mathcal{R}$ for a given event trigger $t$, with $\mathcal{R}$ being the role type ontology.
Given a target task (such as NER) test sample $x_{test}$ and k demonstrations ${(x_{i},y_{i})}_{i=1}^{k}$, we use a frozen large language model to generate answer $y_{test}$ auto-regressively. Our goal is to retrieve $k$ demonstrations from the candidate pool $P$ that most closely match $y_{test}$ to $y_{truth}$. It is worth noting that our candidate pool $P$ contains a variety of information extraction tasks, such as NER, RE, ED and EAE.
Figure 2: The overall architecture of RUIE. The training process consists of three steps: 1) the sparse retriever bm25 initializes a candidate set, which is then scored by the LLM. 2) a keyword-enhanced reward model captures fine-grained information. 3) a bi-encoder dense retriever is trained using contrastive learning and knowledge distillation. During inference, the trained dense retriever selects the best demonstrations from candidate pool $P$, and passes them to the LLM to produce the output.
3.2 LLM Preference Scoring
Sample Format. In the candidate pool $P$, each task sample $(x_{i},y_{i})$ consists of four parts: (1) Task Name: name of a specific IE task, such as “Named Entity Recognition”. (2) Schema: task ontology presented in the form of python list. (3) Input: input context to be extracted. (4) Output: structured output linearized by natural language, such as “Entitytype1: EntityName1; …”.
Score examples using LLM. Previous retrieval based information extraction methods usually only consider the text similarity between the input text and the candidate pool examples, ignoring the preference of the LLM. We assume that the LLM knows what good examples are, specifically, good examples can maximize the probability that the model outputs ground-truth $y_{truth}$. Therefore, given an input $s=(x,y)$ in the training set, we enumerate each example $e_{i}=(x_{i},y_{i})$ in the candidate pool $P$. To keep with the inference stage, we concatenate the extraction instruction $I$ (see Appendix D), example $e_{i}$, and input $s$ into the model. We use the token-level average log-likelihood of $y$ as the score between input $s$ and example $e_{i}$:
$Score(s,e_{i})=log~{}p(y|I;x_{i};y_{i};x)$ (1)
Then we rank the all candidate examples in descending order and select top $k$ and last $n$ examples as positive and negative examples respectively.
Initialize candidates. In practice, the number of examples in the candidate pool $P$ is usually very large. To reduce the costs of LLM scoring, given an input sample $s=(x,y)$, we use a sparse retriever bm25 Robertson and Zaragoza (2009) to retrieve top-k candidate samples $\{(x_{i},y_{i})\}_{i}^{k}$ in candidate pool $P$ as the set of scoring candidates. Choosing an appropriate k value can greatly improve the scoring efficiency and ensure that the candidate set contains both positive and negative samples.
3.3 Keyword-enhanced Reward
Some studies have shown that the alignment of query and candidate examples on fine-grained information such as entities and relations is more important than the alignment on coarse-grained information, like sentence semantics. Therefore, in order to make the fine-grained information between the input query and the candidate examples fully interactive, we propose a keyword-enhanced training strategy based on cross-encoder. On the one hand, the strategy aligns the fine-grained information through keyword, and on the other hand, the cross-encoder realizes the full interaction between query and candidate examples by accessing the ground-truth. Specifically, given a training sample $(x,y)$, for each information snippets $(s,o)$ in $y$, where $o$ is the label of span $s$, we add a pair of special tags “<Keyword>” and “</Keyword>” around $s$ in context $x$ (Figure 2). Given an enhanced input $(x^{\prime},y^{\prime})$, we sample a positive example $(x^{\prime}_{+},y^{\prime}_{+})$ in the top-k of the enhanced ranked candidates and take the last-n as the negative examples $(x^{\prime}_{-_{i}},y^{\prime}_{-_{i}})$. We train the cross-encoder using the cross-entropy loss:
$\mathcal{L}_{reward}=-log\frac{e^{s(x^{\prime},y^{\prime},x^{\prime}_{+},y^{\prime}_{+})}}{e^{s(x^{\prime},y^{\prime},x^{\prime}_{+},y^{\prime}_{+})}+\sum_{i=1}^{n}{e^{s(x^{\prime},y^{\prime},x^{\prime}_{-_{i}},y^{\prime}_{-_{i}})}}}$ (2)
3.4 UIE Retriever Training
In the test phase, we have no access to the ground-truth output $y$ corresponding to the input $x$. So we cannot directly use the above cross-encoder as the retriever. To improve retrieval efficiency, we build a UIE retriever based on a bi-encoder. Specifically, we encode all samples $e^{\prime}_{i}=(x^{\prime}_{i},y^{\prime}_{i})$ in the keyword-enhanced candidate pool $P$ into vector $h_{e^{\prime}i}$ and build the index in advance. Given an input $x$, we use the same encoder to encode it into a vector $h_{x}$ and calculate matching score between the input $x$ and each example $e^{\prime}_{i}$ using cosine similarity:
$Score(x,e^{\prime}_{i})=\cos{(h_{x},h_{e^{\prime}i})}$ (3)
$h_{e^{\prime}i}$ $=\text{avgpool}(Encoder(x^{\prime}_{i};y^{\prime}_{i}))$
$h_{x}$ $=\text{avgpool}(Encoder(x))$
where “avgpool” refers to average pooling and “;” refers to concatenation operation.
We use two kinds of supervision signals to train the UIE retriever: (1) In order to make full use of the positive and negative pairs discovered by LLM, we use the Info-NCE loss $L_{contrastive}$ to perform contrastive learning between the positives and in-batch negatives. (2) In order to make full use of the fine-grained alignment information of the reward model, we use the KL divergence to align the output distributions of the reward model and the retriever, calculated as $L_{distill}$, and the final training loss is the weighted sum of the above two losses:
$\mathcal{L}_{retriever}=\mathcal{L}_{distill}+\alpha\mathcal{L}_{contrastive}$ (4)
where $\alpha$ is a hyper-parameter to measure the importance of the above two losses.
Method NER RE ED / EAE
CrossNER FewRel Wiki-ZSL #Avg WikiEvents RAMS CrudeOil News #Avg
Supervised Fine-tuning Methods
UIE 38.37 - - - 5.12 / 1.78 9.25 / 2.14 6.45 / 8.95 6.94 / 4.29
InstructUIE 49.36 39.55 35.20 37.38 11.64 / 5.88 24.27 / 6.21 23.26 / 21.78 19.72 / 11.29
YAYI-UIE 50.39 36.09 41.07 38.58 10.97 / 5.11 18.87 / 8.21 12.45 / 19.74 14.10 / 11.02
LLaMA2-IEPILE 56.50 37.14 36.18 36.66 13.93 / 12.55 23.62 / 11.30 33.87 / 18.47 23.81 / 14.11
Retrieval-based Methods (k-shot=8)
Random 56.61 21.58 23.27 22.42 61.94 / 35.92 19.12 / 22.94 22.16 / 29.73 34.41 / 29.53
E5 45.21 41.50 44.62 43.06 66.09 / 41.18 26.78 / 25.49 46.47 / 53.12 46.45 / 39.93
BGE 49.69 47.30 51.07 49.19 65.40 / 43.12 29.38 / 25.35 48.05 / 53.33 47.61 / 40.60
BM25 59.57 44.27 50.72 47.50 63.79 / 40.11 25.46 / 27.17 54.05 / 53.48 47.77 / 40.25
RUIE 65.41 49.93 53.16 51.55 66.42 / 40.64 34.60 / 26.06 51.50 / 53.79 50.84 / 40.16
RUIE-Deepseek 69.60 57.57 60.14 58.85 71.88 / 44.30 47.72 / 38.51 68.73 / 61.04 62.77 / 47.95
Table 1: Performance (in F1%) on NER, RE, ED and EAE tasks under held-out settings. Bold indicates the highest scores and the second-best scores are underlined. The only difference between RUIE and other retrieval-based methods is that they use different retrievers. It’s worth noting that the model size of InstuctUIE and LLaMA2-IEPILE are 11B and 13B, respectively, while the default model size of RUIE is 8B.
4 Experiment Setup
4.1 Datasets
To exhaustively evaluate the generalization ability of RUIE, we collect 31 held-in and 8 held-out datasets. We used the training set of the held-in and held-out datasets to form the candidate pool $P$. We constructed the retriever training set based on the held-in dataset, specifically, we sampled 10000 samples from each dataset and included all examples from datasets with less than 10000 samples. The Held-out dataset samples was completely unseen during training, and the UIE Retriever was tested on the test set of held-in and held-out data after training.
4.2 Metrics
We employ span-based Micro-F1 to evaluate our method performance. For NER, an entity is considered correct if the entity span and type are correctly predicted. For RE, a relation is considered correct if relation type, subject entity, and object entity match the golden annotation. For EE task, we report two evaluation metrics: (1) ED: an event trigger is correct if the event type and the trigger word are correctly predicted. (2) EAE: an event argument is correct if its role type and event type match a reference argument mention.
4.3 Baseline Methods
We select three kinds of models for comparative analysis: (1) Traditional UIE method Lu et al. (2022), which is based on relatively small language models (i.e., million-level parameters). (2) Instruction-tuning based methods: Wang et al. (2023b) converts many public IE datasets to instruction-tuning datasets using natural language. Gui et al. (2024); Xiao et al. (2024b) constructs large instruction-tuning datasets with json output format. Li et al. (2024) uses code format to model IE structure. (3) Retrival-based methods: The first is sparse retriever bm25Robertson and Zaragoza (2009). The others are dense retriever Wang et al. (2024a); Xiao et al. (2024a) which are trained in a contrastive manner with large-scale text pair dataset. It is worth noting that Wang et al. (2023a); Wan et al. (2023) presuppose the entity span in the sentence, which is inconsistent with our experimental settings, so we did not compare them.
4.4 Implementation Details
Unless otherwise stated, we use llama3 as the scoring model and llama3.1-instruct for inference by default. For each test sample, we retrieve 8 examples for inference by default. We trained the keyword enhanced reward model and retriever based on ELECTRA-base and E5-base respectively. More detailed are in Appendix B:
5 Results and Analyses
5.1 Generalization to Held-out Tasks
In the held-out tasks generalization experiment, SFT-based methods are directly tested on held-out tasks after training on held-in tasks. After training on held-in tasks, RUIE retrieves demonstrations from held-out tasks for LLM. The remaining retrieval-based methods retrieve demonstrations for LLM directly at held-out tasks. Results are shown in Table 1.
RUIE achieves the best performance in four information extraction tasks, and demonstrates rapid generalization to new tasks. Compared with SFT-based methods, the generalization ability of RUIE is significantly better. The performance of SFT-based methods decreases seriously on datasets not seen during training, with the drop becoming more pronounced as task difficulty increases. For instance, in the ET and EAE tasks, the best-performing LLaMA2-IEPILE model only achieves F1 scores of 23.81 and 14.11, making it impractical for real-world use. In contrast, RUIE delivers improvements of 8.91, 14.89, 27.03, and 26.05 on NER, RE, ET, and EAE tasks, respectively, despite using smaller model.
Compared to general-purpose retrievers, RUIE retrieves higher-quality examples. BM25 is a strong baseline and outperforms semantic similarity based retrieval methods on average, confirming the alignment of fine-grained information is more important in IE example retrieval. By explicitly modeling fine-grained information, RUIE introduces LLMs preference as a supervision signal during retrieval training. Compared to BM25, which has the best overall performance, RUIE achieves 5.84, 4.05, and 3.07 improvements on NER, RE, and EET IE tasks respectively.
However, RUIE does not achieve the performance improvement on EAE. We believe there are two main reasons. First, compared to other IE tasks, event extraction data is more scarce, leading to insufficient training during the pre-training and alignment stages. Additionally, EAE requires the model to extract event parameters while understanding the event structure, making it more challenging than other tasks. Furthermore, the amount of EAE training data is less than other subtasks, limiting RUIE’s generalization performance on new EAE tasks.
Dataset SOTA E5 BM25 RUIE
[…]
ACE 2005 $\bigstar$ 86.66 42.53 48.35 55.86
AnatEM $\bigstar$ 90.89 41.71 46.85 51.58
bc2gm $\bigstar$ 85.16 43.48 45.99 49.78
bc4chemd $\spadesuit$ 90.56 48.29 50.39 55.03
bc5cdr $\bigstar$ 89.59 71.41 72.76 74.49
Broad Tweet $\blacktriangle$ 83.52 61.92 56.83 69.25
CoNLL 2003 $\blacktriangle$ 96.77 68.09 68.50 78.34
FabNER $\clubsuit$ 82.90 35.24 39.61 38.51
FindVehicle $\spadesuit$ 99.45 68.86 73.53 92.26
GENIA_NER $\spadesuit$ 78.29 53.23 57.64 59.85
HarveyNER $\bigstar$ 88.79 29.74 33.59 37.72
MultiNERD $\clubsuit$ 96.10 82.50 81.71 88.97
ncbi $\bigstar$ 90.23 54.60 59.14 58.81
Ontonotes $\bigstar$ 90.19 48.31 49.85 62.94
PolyglotNER $\blacktriangle$ 70.85 49.15 49.21 53.43
TweetNER7 $\blacktriangle$ 66.99 49.93 49.79 53.17
WikiANN en $\clubsuit$ 87.00 64.07 58.27 68.81
WikiNeural $\bigstar$ 91.36 74.58 73.28 81.35
Avg 86.99 54.23 55.99 62.46
[…]
Dataset SOTA E5 BM25 RUIE
ADE corpus $\spadesuit$ 82.31 70.39 65.44 71.24
Conll04 $\bigstar$ 78.48 41.42 44.52 54.61
GIDS $\bigstar$ 81.98 27.49 30.44 40.03
Kbp37 $\clubsuit$ 78.00 12.76 11.79 19.78
NYT $\spadesuit$ 94.04 55.95 74.85 72.30
NYT11 $\blacktriangle$ 57.53 31.16 33.49 40.92
SciERC $\spadesuit$ 45.89 12.25 16.62 20.36
Semeval RE $\bigstar$ 73.23 23.07 21.84 36.77
Avg 73.93 34.31 37.37 44.50
Table 3: Performance (in F1%) comparison on RE tasks under held-in setting
Dataset SOTA E5 BM25 RUIE
ACE2005 $\bigstar$ 77.13 42.90 39.06 53.41
CASIE $\bigstar$ 67.80 29.35 39.69 40.46
PHEE $\bigstar$ 70.14 37.18 59.21 47.16
Avg 71.69 36.48 45.99 47.01
Table 4: Performance (in F1%) comparison on ED tasks under held-in setting.
Dataset SOTA E5 BM25 RUIE
ACE2005 $\bigstar$ 72.94 39.06 38.29 44.29
CASIE $\blacktriangle$ 64.23 39.69 40.85 43.76
PHEE $\blacktriangle$ 77.19 59.21 59.10 65.13
Avg 71.45 45.99 46.08 51.06
Table 5: Performance (in F1%) comparison on EAE tasks under held-in setting.
5.2 Results on Held-in Tasks
In the held-in tasks experiment, both supervised fine-tuning methods and RUIE are trained on held-in tasks and tested on their corresponding test sets. The remaining retrieving-based methods directly retrieve k examples from held-in tasks for LLM. SFT-based methods include Lu et al. (2022); Wang et al. (2023b); Gui et al. (2024); Xiao et al. (2024b); Li et al. (2024). As some baselines only report results for specific IE tasks, we report the SOTA results of the above methods in each dataset, denoted as “SOTA” in the tables.
The results of NER, RE, ED, and EAE are shown in Tables 2, 3, 4, and 5. We conclude two main findings: (1) SFT-basd methods provide LLMs with a comprehensive understanding of IE through large-scale IE datasets training, achieving superior results on held-in tasks compared to the RUIE, which utilizes a frozen LLM. However, RUIE achieved an F1 score of 92.26 on the FindVehicle, showing the strong potential of retrieval-based UIE. (2) RUIE showed markedly superior to generic retrievers, with improvements of 6.47, 7.13, 1.35, and 4.98 on the NER, RE, ED, and EAE tasks, respectively. This proves the effectiveness of introducing large model preference and keyword enhancement during the retriever training process.
LLMs NER RE ED / EAE
RUIE 68.24 51.55 50.84 / 40.16
- keyword-enhanced 68.64 51.37 48.82 / 39.47
- reward model 63.69 32.24 43.57 / 35.83
- distill loss 62.78 33.73 39.77 / 31.58
Table 6: Ablation study across all held-out datasets. “- keyword-enhanced” trains the reward model and retriever without keyword enhancement. “- reward model” trains the retriever with distill loss on LLMs scores and contrastive loss on ranked candidates. “- distill loss” trains the retriever soly with contrastive loss on ranked candidates.
5.3 Ablation Study
We conduct an ablation study across all held-out datasets to underscore the effectiveness of the key innovations in our work (Table 6). Removing the keyword enhancement led to a 0.62 decrease in the average F1-score across four tasks, highlighting the value of fine-grained information alignment during retriever training. After removing distill loss, the average f1-score of the four tasks decreases by 10.73, which is because the rank of candidate examples only reflects the relative distribution of LLs preferences. However, knowledge distillation based on KL-divergence can align the absolute distribution of LLMs preferences, which confirms the importance of using LLMs preferences as supervision signals. After removing the reward model, the average f1-score of the four tasks decreases by 8.87, indicating the log-likelihood from LLMs is not suitable as a direct knowledge source for distillation. Because the average log-likelihood is not a true probability distribution, and its values tend to cluster within a narrow range, making it less effective as a target distribution within the KL-divergence framework.
5.4 The Effects of Different K-shot Demonstrations
Figure 3: Performance (in F1%) comparison by varying k-shot demonstrations.
We use the default experimental setup to investigate the impact of varying k-shot demonstrations. Experiments were conducted across all held-in and held-out datasets, and we reported the average F1 scores for each task across all datasets, as depicted in Figure 3. We summarized three findings: (1) The performance across all tasks improved with an increase in k, with the most notable enhancement occurring as k rose from 0 to 2, indicating that few-shot demonstrations are crucial for LLMs to tackle tasks effectively. (2) The task performance does not consistently improve with the increment of k. For instance, there is a performance decline when k increased from 12 to 16 in the RE task. We surmise that one contributing factor is the limitation of the model’s context length, while another is the additional noise introduced by an excess of examples, which can adversely affect model performance. Therefore, we set k to 8, which is able to take into account both inference efficiency and extraction performance. (3) Model performance is influenced by the complexity of the task. NER has more data and is easier compared to other information extraction tasks, exhibited significantly better zero-shot performance than other tasks, while LLM was virtually incapable of completing the RE and ED tasks in a zero-shot scenario. Even as the number of examples k increased, the model’s performance on NER remained superior to that of other information extraction tasks.
5.5 The Effects of Different Scoring LLMs
LLMs NER RE ED / EAE
GPTNeo 68.27 50.78 47.86 / 40.35
LLaMA3-instruct 68.14 51.24 49.07 / 40.68
LLaMA3 68.24 51.55 50.84 / 40.16
Table 7: Performance (in F1%) comparison of different scoring LLMs on NER, RE, ED, EAE tasks across all held-out datasets.
To investigate the impact of various scoring LLMs on the extraction performance, we conducted experiments across all held-out datasets, with results presented in Table 7. We have two findings: 1) The size of the scoring LLM has a minor influence on the final performance. Although the LLaMA3 (8B) outperformed the GPTNeo (2.7B) by 2.98 in the ED task, their performance was comparable in NER, RE, and EAE tasks. This suggests that models of different sizes exhibit similar preferences for certain tasks, allowing one to select an appropriately sized LLM based on computational resources. 2) The base version of the LLM is more suitable as a scoring model than the instruct version. We calculated the mean and variance of the scores for the positive samples (top-3) and the negative samples (last-16) for both versions of the LLM (Table 10). Although the mean difference between positive and negative sample scores of instruct version LLM is larger, the variance of positive and negative sample scores is significantly larger than that of base version LLM, which introduces instability to the subsequent training of the reward model, resulting in the final extraction performance inferior to base version.
5.6 The Effects of Different Inference LLMs
Table 8 shows the performance of various reasoning LLMs across all held-out datasets. We find the ability of the reasoning LLMs significantly influence the extraction performance, which is expressed in two aspects: 1) Model size: For the same Qwen1.5, an increase in model size from 7B to 14B resulted in an average performance enhancement of 9.21 across four tasks. 2) Model type: Within the LLaMA series, LLaMA3.1, which is an advancement over LLaMA3, achieved an average performance improvement of 0.87 across four tasks. Since all the tasks are in English, the LLaMA series models consistently outperformed the Qwen series. Furthermore, RUIE can effectively adapt to both locally deployed models and API-based models, with Deepseek demonstrating the most superior performance.
LLMs NER RE ED / EAE
Qwen1.5-7b 54.70 33.86 33.84 / 26.90
Qwen1.5-14b 63.35 44.53 40.09 / 32.23
LLaMA3-8b 67.19 53.65 46.83 / 39.64
LLaMA3.1-8b 68.24 51.55 50.84 / 40.16
Deepseek 72.24 58.85 62.77 / 47.95
Table 8: Performance (in F1%) comparison of different inference LLMs on NER, RE, ED, EAE tasks across all held-out datasets. Deepseek refers to the deepseek-chat-v2. All other models are instruct variant, which are further fine-tuned on instruction-tuning datasets after pre-training.
6 Conclusion
In this paper, we propose RUIE, the first trainable retrieval-based framework designed for the UIE, which reduces the computational cost of fine-tuning and enables fast generalization to unseen tasks. We collected 31 datasets to train RUIE and tested it on 8 unseen tasks. The experimental results show that the introduction of LLMs preference and keyword-enhanced reward in RUIE effectively improves the information extraction ability of LLMs.
Limitations
Context length limitation: RUIE currently focuses on sentence-level UIE. On the one hand, it is difficult to meet the retrieval requirements of long documents due to the length limitation of the retriever is 512. On the other hand, due to the length limitation of LLMs, k-shot cannot be increased in the case of long documents, which limits the performance of the model.
Gap between RUIE and the SFT-based methods in seen tasks: The SFT-based methods inject UIE-specific knowledge into the LLMs through fine-tuning, while RUIE frozens the LLMs and can only leverage the knowledge from the pre-training phase. How to better stimulate the information extraction ability of LLMs under controllable cost is a worthy study direction.
English corpus only: RUIE is currently only trained and tested on English data. In the future, we hope to expand RUIE to more languages.
Appendix A Data Details
Our candidate pool $P$ mainly consists of IE INSTRUCTIONS Wang et al. (2023b) and IEPILE Gui et al. (2024). There are 22 NER datasets: ACE2004 Mitchell et al. (2005), ACE2005 Walker and Consortium (2005), Broad Twitter Derczynski et al. (2016), CoNLL2003 Sang and Meulder (2003), MultiNERD Tedeschi and Navigli (2022), Ontonotes Hovy et al. (2006), Polyglot-NER Al-Rfou et al. (2014), tweetNER7 Ushio et al. (2022), wikiANN Pan et al. (2017), wikineural Tedeschi et al. (2021), AnatEM openbiocorpora (2015), bc2gm Kocaman and Talby (2020), bc4chemd Krallinger et al. (2015), bc5cdr Li et al. (2016), FabNERKumar and Starly (2021), FindVehicle Guan et al. (2023), GENIA Kim et al. (2003b), HarveyNER Chen et al. (2022), MIT Movie Liu et al. (2019) MIT Restaurant Liu et al. (2019) ncbi-disease Dogan et al. (2014). For the RE task, we utilize 10 datasets including ADE corpus Gurulingappa et al. (2012), CoNLL04 Roth and tau Yih (2004), GIDS Jat et al. (2018), kbp37 Zhang and Wang (2015), NYT Riedel et al. (2010), NYT11 HRL Takanobu et al. (2018), SciERC Luan et al. (2018), semeval RE Hendrickx et al. (2010), FewRel Han et al. (2018) and Wiki-ZSL Chen and Li (2021). For the EE task, ACE2005 Walker and Consortium (2005), CASIELu et al. (2021), GENIAKim et al. (2003a), PHEESun et al. (2022), CrudeOilNews Lee et al. (2022), RAMS Ebner et al. (2020) and WikiEvents Li et al. (2021) are used.
Appendix B Implementation Details
Reward Retriever
learning rate 1e-5 3e-5
batch size 64 128
training steps 3000 3000
$\alpha$ 0.2
positives top-3 top-3
negatives last-16 in-batch negatives
input length 512 512
Table 9: Hyperparameters for reward and retriever model training.
We finished LLM scoring, reward and retriever training on two 3090 GPUs. In order to balance the efficiency and performance of the scoring process, we used bm25 to initialize 100 candidates for each sample in the training set. Due to the context length limitation of the LLMs, we retrieve 8 samples for each query and set the maximum input length to 1792 and the maximum generation length to 256. We performed inference using vllm Kwon et al. (2023) on a single 3090 GPU. In order to ensure the reproduction of the results, we use a greedy decoding strategy and set the temperature to 0.
Appendix C Score Mean and Variance of different LLMs
Model Positives Negatives
mean variance mean variance
base -0.39 0.18 -1.02 0.93
instruct -0.36 0.35 -1.23 2.36
Table 10: Score Mean and Variance of different LLMs. “base” indicates LLaMA3-8B and “instruct” indicates LLaMA3-instruct-8B. We reported score mean and variance in positives and negatives separately.
Task Dataset #Schema #Train #Test Training Evaluation
NER ACE2004 7 6202 812 $\surd$ $\surd$
ACE2005 7 7299 1060 $\surd$ $\surd$
Broad Tweet 3 5334 2001 $\surd$ $\surd$
CoNLL2003 4 14041 3453 $\surd$ $\surd$
multiNERD 16 134144 10000 $\surd$ $\surd$
Ontonotes 18 59924 8262 $\surd$ $\surd$
Polyglot-NER 3 393982 10000 $\surd$ $\surd$
tweetNER7 7 7111 576 $\surd$ $\surd$
Wikiann 3 20000 10000 $\surd$ $\surd$
wikineural 3 92729 11597 $\surd$ $\surd$
anatEM 1 5861 3830 $\surd$ $\surd$
Bc2gm 1 12500 5000 $\surd$ $\surd$
Bc4chemd 1 30682 26364 $\surd$ $\surd$
Bc5cd 2 4560 4797 $\surd$ $\surd$
FabNER 12 9435 2064 $\surd$ $\surd$
FindVehicle 21 21565 20777 $\surd$ $\surd$
GENIA 5 15023 1854 $\surd$ $\surd$
HarveyNER 4 3967 1303 $\surd$ $\surd$
Ncbi-disease 1 5432 940 $\surd$ $\surd$
[…]
RE ADE corpus 1 3417 428 $\surd$ $\surd$
CoNLL2004 5 922 288 $\surd$ $\surd$
GIDS 4 8526 4307 $\surd$ $\surd$
Kbp37 18 15917 3405 $\surd$ $\surd$
NYT 24 56196 5000 $\surd$ $\surd$
NYT11 HRL 12 62648 369 $\surd$ $\surd$
SciERC 7 1366 397 $\surd$ $\surd$
Semeval RE 10 6507 2717 $\surd$ $\surd$
[…]
EE ACE2005 33(22) 3342 293 $\surd$ $\surd$
CASIE 5(26) 3751 1500 $\surd$ $\surd$
GENIA 5(0) 15023 1854 $\surd$ $\surd$
PHEE 2(16) 2898 968 $\surd$ $\surd$
[…]
Table 11: Statistical data of all 39 IE datasets. Datasets with $\dagger$ are held-out datasets, which is unseen during retriever training stage. CrossNER Liu et al. (2021) is divided into five subsets for our statistical analysis. For the held-out datasets, we only report their test set.
Appendix D Instruction Format
Detailed instruction format is listed in Table 12.
Task Instruction
NER Please analyze the given schema and extract all named entities from the provided input. Follow these instructions carefully:
1. Output Format: Present the extracted entities in this structured format: “EntityType1: EntityName1; EntityType2: EntityName2; …”.
2. Include Only Present Entities: Only output entities that actually exist in the input. Ignore any entities that are not mentioned.
3. No Entities Response: If the input contains no named entities, respond with “None”.
4. Examples Handling: If examples are provided for reference, use them to understand the annotation criteria, but do not extract entities from these examples.
RE Please analyze the given sentence and extract subjects and objects that have a specific relation, according to the provided schema. Follow these guidelines:
1. Output Format: Format your output as follows: “relation1: subject1, object1; relation2: subject2, object2; …”.
2. Include Only Present Entities: Only output entities that actually exist in the input. Ignore any entities that are not mentioned.
3. No Relations Response: If the sentence contains no relations, respond with “None”.
4. Example Usage: If examples are provided, use them to understand the annotation criteria. Do not extract relations from these examples.
ED Please analyze the given schema and extract all event triggers from the provided input. Follow these instructions carefully:
1. Output Format: Present the extracted triggers in this structured format: “EventType1: TriggerName1; EventType2: TriggerName2; …”.
2. Include Only Present Triggers: Only output triggers that actually exist in the input. Ignore any triggers that are not mentioned.
3. No Triggers Response: If the input contains no event triggers, respond with “None”.
4. Examples Handling: If examples are provided for reference, use them to understand the annotation criteria, but do not extract triggers from these examples.
EAE Please extract and list arguments of specified types for a given event type and trigger, according to the provided schema. Follow these instructions carefully:
1. Output Format: Format your output as follows: “ArgumentType1: argument1; ArgumentType2: argument2; …”.
2. Include Only Present Arguments: Only include event arguments that are present in the input sentence. Ignore any event arguments that are not mentioned.
3. No Arguments Response: If the sentence contains no event arguments, respond with “None”.
4. Example Usage: If examples are provided, use them to understand the annotation criteria. Do not extract event arguments from these examples.
Table 12: Detailed instruction format across four IE tasks.